Exercise 13-11 Profitability analysis LO P3

Simon Company’s year-end balance sheets follow.

At December 31 2017 2016 2015
Assets
Cash $ 31,400 $ 34,700 $ 36,200
Accounts receivable, net 88,500 64,200 58,600
Merchandise inventory 107,339 84,200 50,300
Prepaid expenses 11,047 9,166 3,628
Plant assets, net

331,714

257,734 191,272
Total assets $ 570,000 $ 450,000 $ 340,000
Liabilities and Equity
Accounts payable $ 143,349 $ 76,050 $ 44,880
Long-term notes payable secured by
mortgages on plant assets
106,088 105,570 73,637
Common stock, $10 par value 162,500 162,500 162,500
Retained earnings 158,063 105,880 58,983
Total liabilities and equity $ 570,000 $ 450,000 $ 340,000


The company’s income statements for the years ended December 31, 2017 and 2016, follow.

For Year Ended December 31 2017 2016
Sales $ 741,000 $ 535,500
Cost of goods sold $ 452,010 $ 348,075
Other operating expenses 229,710 135,482
Interest expense 12,597 12,317
Income taxes 9,633 8,033
Total costs and expenses 703,950 503,907
Net income $ 37,050 $ 31,593
Earnings per share $ 2.28 $ 1.94


Additional information about the company follows.

Common stock market price, December 31, 2017 $ 28.00
Common stock market price, December 31, 2016 26.00
Annual cash dividends per share in 2017 0.22
Annual cash dividends per share in 2016 0.11


To help evaluate the company's profitability, compute the following ratios for 2017 and 2016:

1. Return on common stockholders' equity.
2. Price-earnings ratio on December 31.
3. Dividend yield.

Compute the return on common stockholders' equity for each year.

Return On Common Stockholders’ Equity
Choose Numerator: / Choose Denominator: = Return On Common Stockholders’ Equity
Net income Preferred dividends / Average common stockholders’ equity = Return on common stockholders’ equity
2017 $37,050 0 / $162,500 = 22.8 %
2016 $31,593 0 / = 0 %

Compute the price-earnings ratio for each year. (Round your answers to 2 decimal places.)

  • Price-Earnings Ratio
    Choose Numerator: / Choose Denominator: = Price-Earnings Ratio
    Market price per common share / Earnings per share = Price-earnings ratio
    2017 / = 0
    2016 / = 0

Compute the dividend yield for each year. (Round your answers to 2 decimal places.)

Dividend Yield
Choose Numerator: / Choose Denominator: = Dividend Yield
Annual cash dividends per share / Market price per share = Dividend yield
2017 / = 0 %
2016 / = 0 %
Answer #1

(1)

Return on common stockholders' equity = (net income - preferred dividends)/Average common stockholders’ equity

for 2017:

Average common stockholders’ equity = {($162500 + $158063) + ($162500 + $105880)}/2 = $294471.5

Return on common stockholders' equity = ($37050 - 0)/$294471.5 = 12.58%

for 2016:

Average common stockholders’ equity = {($162500 + $105880) + ($162500 + 58983)}/2 = $244931.5

Return on common stockholders' equity = ($31593 - 0)/$244931.5 = 12.90%

(2)

Price-earnings ratio = market price per share/earning price per share

for 2017

price earning ratio = $28/$2.28 = 12.28 times

for 2016:

price earning ratio = $26/$1.94 = 13.40 times

(3)

Dividend yield = dividend per share/market price per share

for 2017:

dividend yield = $0.22/$28 = 0.79%

for 2016:

dividend yield = $0.11/$26 = 0.42%